OpenStudy (anonymous):

what is the slope of the line, 3x-2y=8?

OpenStudy (anonymous):

bring it in standard y=mx +c form

OpenStudy (anonymous):

the value of m is the slope

OpenStudy (anonymous):

i am getting 8/5 as the answer. i wanna know if that's right.

OpenStudy (anonymous):

it is -3/2

OpenStudy (anonymous):

@ vbkotecha, can u solve it step by step here, so that i know where i went wrong? Thanks a tonne though :)

OpenStudy (anonymous):

3x-2y=8 So bring 2y to another side, it becomes 2y=3x-8 now divide the whole equation by 2 therefore it becomes y=3x/2 -4 the coefficient of x is the slope.And there you have your answer. Sorry I had posted -3/2.......the answer as I figure out...youve got it!

OpenStudy (anonymous):

Well,don't mention it(your thanking me)

OpenStudy (anonymous):

@vbkotecha :), i have no idea how i got 8/5, messed it up like real bad :)

OpenStudy (anonymous):

positive 3/2

OpenStudy (anonymous):

its not -3/2. its +3/2. bringing 3x to the other side means making the sign opposite in this case making it minus -3/2. And then when you divide it by -2 this makes the equation positive.

OpenStudy (anonymous):

I already mentioned my mistake before 5 replies man!! I tell again.....its +3/2 GOD!!!!

OpenStudy (anonymous):

OH LOL. My fault, I skimmed through the replies and I thought you were showing how you did the problem. Sorry again xD

OpenStudy (anonymous):

Oh its alright...sure thanks!

OpenStudy (anonymous):

3x-2y=8 2y=3x-8 y=3/2x-8/2 so slope is +3/2

OpenStudy (anonymous):

Lol, you tell again? vbkotecha. Of the two post you put the answer on, both show -3/2. You said "Sorry I had posted -3/2.......the answer as I figure out...youve got it!," makes no sense. Correct yourself next time if u make an error.